OpenStudy (anonymous):

Solve for t. c=rt+k 1. t=c-k/r 2. t=c+k/r 3. t=c/r-k 4. t=c-r/k

OpenStudy (alexandervonhumboldt2):

subtract k from both sides: c-k=rt+k-k c-k=rt divide both sides by r rt/r=(c-k)/r t=(c-k)/r

OpenStudy (anonymous):

ok thanks